OpenStudy (anonymous):

What is an equation of a line (in point-slope form) that passes through the points (-3,-2) and (2,3)?

OpenStudy (anonymous):

y+2=(3+2)/(2+3) *(x+3) y+2=(x+3) x-y=-1

OpenStudy (anonymous):

point slope form is (y-y1)=m(x-x1)